On friday the snow started. It has not been a very heavy snowfall, but has been pretty continuous. Day 2 was snowed out and we expect Day 3 to be lost as well.
In the meantime, Vanessa aligned the retroreflector on the left side and tested the AO system. She gave it all sorts of abuse (large tilts, simulated turbulence, calling it names) but the system showed no sign of the failure we saw in November. This is good news,and we are confident the left AO system is ready to go.
The AO GUI showing a big tilt on the mirror, right before the safety system takes over. This worked as expected.
Denis, Elwood, Andy, and Phil worked on realignment of the beamcombiner in the afternoon and evening. We slightly adjusted the ellipse on the left side to better align the beamcombiner to the telescope. A loose optic was identified and fixed on the right side. The beamcombiner artificial source was then able to be aligned.
In the evening we scanned (and scanned) for fringes. The optics have to have equal pathlength to within 50 microns and this took us awhile to find. Later in the night we found it. We then were able to successfully close the pahtlength loop. This milestone represents the first end-to-end closed loop verification of the LBTI system. Elwood and Denis worked on refining the loop parameters. In open loop the system showed >600 nm RMS pathlength variation. In closed loop we were able to consistently reduce it to 300 nm.
Plot of the pathlength RMS over a couple minutes. The "bump" in the center is when the loop is open. Times 0-70s and 100-130 s are when the loop is closed.